# Squeezing and Entanglement Dynamics in Phase-Sensitive Non-Hermitian Systems

### v1

@@ Abstract
Introduction.—Although any Hermitian operator has real eigenvalues, hermiticity is not a necessary condition to obtain real eigenvalues. Recent studies have shown that any Hamiltonian satisfying parity-time (PT) symmetry ($[H,PT]=0$) can still possess real eigenvalues [1, 2, 3]. Such systems undergo a phase transition during which the PT symmetry is spontaneously broken. The eigenvalues of the Hamiltonian transition from real to complex at a specific point in the parameter space are known as the exceptional point (EP). At the EP, both the eigenvalues and eigenvectors of the Hamiltonian coalesce, leading to a range of novel phenomena [4, 5, 6, 7, 8, 9, 10]. Corresponding to PT symmetry, anti-PT (APT) symmetry, as another form of non-Hermitian symmetry ($\{H,PT\}=0$), also undergoes spontaneous APT symmetry breaking during the APT phase transition, accompanied by the emergence of EPs [11, 12].
@@
Over the past decade, PT symmetry [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and APT symmetry [26, 27, 28, 29, 30, 31, 32, 33] in various physical systems such as photonics, acoustics, and ultracold atoms have been extensively explored, leading to significant experimental and theoretical advancements. To realize PT symmetry in a physical system, the system typically exhibits linear gain and loss [3, 4, 26, 34]. However, these optical gain and loss mechanisms present significant challenges for achieving PT symmetry in the purely quantum realm [35] because the associated Langevin noise disrupts PT symmetry in quantum systems [36, 37]. Currently, there are two approaches to circumvent this issue: one is a passive scheme based on loss modulation, and the other is the Hamiltonian dilation method [37, 38, 39, 40, 41, 42]. By contrast, APT symmetry can be realized without gain and loss [29, 43, 44, 45]. Recent theoretical research indicates that APT symmetry can be realized in a two-mode bosonic system without involving linear gain and loss of classical fields. This suggests the potential for achieving quantum APT symmetry free from Langevin noise, which can be implemented in nonlinear wave-mixing processes [45].
Recent studies focusing on quantum entanglement [46, 47] have demonstrated the unique properties of PT symmetry and APT symmetry in both two-qubit systems [48] and linear optical systems [49]. However, physical systems that simultaneously exhibit both PT symmetry and APT symmetry have not yet been explored.
@@
In this study, we construct a phase-sensitive non-Hermitian wave mixing model that not only possesses inherent single-mode APT symmetry but also achieves two-mode quadrature-PT symmetry by modulating the phase of the pump field. This aspect has not been explored in previous studies. Classical-nonclassical polarity (CNP) reveals the unique impact of APT symmetry on the nonclassical properties of a system. Specifically, significant changes in the single-mode squeezing dynamics and two-body Gaussian entanglement dynamics of the system are observed as it traverses EP. Furthermore, the results from the inseparability criterion provide deeper insights into the relationship between two-mode quadrature-entanglement and the phase of the pump field. They reveal the influence of single-mode APT symmetry and two-mode quadrature-PT symmetry on the entanglement. Specifically, when the phase of the pump field satisfies $\cos\phi=0$, the system simultaneously exhibits single-mode APT symmetry and two-mode quadrature-PT symmetry. Under these conditions, the two-mode quadrature-entanglement undergoes sudden death in the two-mode quadrature-PT-broken region and EP, while exhibiting periodic revival characteristics in the two-mode quadrature-PT-symmetric region. Conversely, when the phase of the pump field meets the condition $\cos\phi=\pm 1$, the phase transition of the system from the APT-symmetric region to the APT-broken region is accompanied by the emergence of quantum entanglement from nonexistence (symbolizing a transition from classical to quantum). Our work further investigates the unique non-Hermitian properties of nonlinear systems, exploring the relationship between two-mode quantum entanglement and the phase of the pump field. This deeper examination reveals the entanglement generation and evolution process in the corresponding nonlinear system of the model, offering new perspectives for quantum information processing.
@@
Figure 1: (a) A non-Hermitian system composed of a four-wave mixing process and a phase-controllable pump source. $\chi$ represents the nonlinear susceptibility. $\varphi$ is the phase of the pump field, $\theta$ is the angle between the pump field and the signal (idler) field, and $L$ represents the total length of the nonlinear medium along the optical axis of the pump field. (b) APT phase transition process of the eigenvalues $\lambda_{\pm}$.
@@
Classical-Nonclassical Polarity.—To investigate the intrinsic connection between non-classical properties and the non-Hermitian nature of this phase-sensitive model, we first study Classical-Nonclassical Polarity (CNP) [50] in our model.
Using the covariance matrix $\gamma_{si}$ (see supplementary materials for more detail), the single-mode and two-mode CNPs can be written as [50]
@@
$P_{1}$ characterizes the single-mode squeezing property of the system. In our model, the signal and idler fields exhibit the same single-mode squeezing properties. $P_{2}$ characterizes the two-mode Gaussian entanglement of the system. When the CNP is greater than 0, it indicates that the system is in a nonclassical state. When the CNP is equal to zero, the system is at the boundary between the classical and nonclassical states. When the CNP is less than 0, the system is in a classical state. The absolute value of the CNP quantifies the distance of the state from the classical-nonclassical boundary.
The nonclassical properties measured by CNP have two important characteristics. First, as shown in Eq. (5a) and (5b), the intensity of the input coherent state is independent of the nonclassical properties measured by the CNP. Second, the phase of the pump field is also independent of the nonclassical properties measured by the CNP, meaning that single-mode squeezing and two-mode Gaussian entanglement are not affected by the phase of the pump field. The nonclassical properties primarily depend on the system parameters $\kappa$ and $\mathrm{\Delta}$, as well as the evolution distance $\kappa L$. Therefore, discussing the CNP connection between single-mode APT symmetry and two-mode quadrature-PT symmetry here does not violate the condition $\cos\phi=0$.
@@
Fig.3 shows the evolution of type-I/type-II entanglement in different phase $\phi$. When the phase $\phi$ satisfies $\cos\phi=0$ (e.g., $\phi=1.5\pi$ or $0.5\pi$), the system simultaneously exhibits single-mode APT symmetry and two-mode quadrature-PT symmetry, as shown by the dashed gray line. By comparing fig.3 ($\mathrm{a}_{1}$), ($b_{1}$), and ($c_{1}$) with fig.3 ($a_{2}$), ($b_{2}$), and ($c_{2}$), it can be seen that the type-I entanglement $E_{1}$ and type-II entanglement $E_{2}$ differ only by a phase of $\pi$. Therefore, the following analysis focuses on the type-I entanglement $E_{1}$. Fig.3($a_{1}$) shows that when the system is in the single-mode APT-symmetric region ($\delta=0.8$) and the phase $\phi$ satisfies $\pi<\phi<2\pi$, the system is still capable of generating entanglement. However, as the system evolves, entanglement is maintained within a very narrow phase range. Outside this range, entanglement briefly exists before rapidly vanishing. When the phase $\phi$ assumes other values (e.g., $0<\phi<\pi$), no entanglement is generated in the system. Fig.3($c_{1}$) shows that in the single-mode APT-broken region ($\delta=1.2$), the evolution of the system entanglement exhibits an oscillatory pattern with a period $T=\kappa\pi/\lambda$. The degree of entanglement can be tuned by varying the phase $\phi$. Interestingly, as shown in Fig.3($b_{1}$), when the system is at the EP ($\delta=1$), by tuning the phase $\phi$ close to 0, entanglement can be maintained over a wider phase range than in the single-mode APT-symmetric region. This demonstrates the significant role of EP control in non-Hermitian systems and phase $\phi$ tuning in the generation of entanglement in this model.
@@
Conclusion.—In summary, unlike previous studies, this study introduces a phase-sensitive non-Hermitian four-wave mixing model. The single-mode effective Hamiltonian of this model satisfies APT symmetry, and by adjusting the phase of the pump field, the two-mode effective Hamiltonian can simultaneously exhibit two-mode quadrature-PT symmetry. By analyzing the system’s non-classical characteristics through CNP and the inseparability criterion, it is found that the single-mode squeezing and two-mode Gaussian entanglement measured by CNP are independent of the phase of the pump field, whereas the two-mode quadrature-entanglement measured by the inseparability criterion is phase-dependent. This is primarily because CNP measures the overall non-classicality of the system, whereas the inseparability criterion assesses the entanglement between quadrature components of the two-mode fields. Consequently, they reveal different aspects of non-classical properties.
Theoretical calculations of the inseparability criterion show that, when the pump phase satisfies $\cos\phi=0$, the system simultaneously exhibits single-mode APT symmetry and two-mode quadrature-PT symmetry. Under these conditions, quantum entanglement undergoes sudden death in the two-mode quadrature-PT-broken region and EP, whereas it exhibits periodic revival characteristics in the two-mode quadrature-PT-symmetric region. When the pump phase satisfies the condition $\cos\phi=\pm 1$, the phase transition of the system from the APT-symmetric region through the EP to the APT-broken region is accompanied by a transition of quantum entanglement from nonexistence to existence, symbolizing a shift from classical to quantum behavior. Our work further explores the unique non-Hermitian properties of nonlinear systems, investigates the relationship between two-mode quantum entanglement and the pump phase, and provides a deeper understanding of the entanglement generation and evolution processes in the model. This study offers new insights into quantum information processing.

### v2

@@ Abstract
Introduction.—While Hermitian operators have real eigenvalues, hermiticity is not required for real eigenvalues. Recent studies indicate that non-Hermitian Hamiltonians with parity-time (PT) symmetry ($[H,PT]=0$) can possess real eigenvalues [1, 2, 3]. These systems experience a phase transition where PT symmetry is spontaneously broken. At a specific parameter point, known as the exceptional point (EP), Hamiltonian eigenvalues shift from real to complex, and eigenvalues and eigenvectors coalesce, leading to novel phenomena [4, 5, 6, 7, 8, 9, 10]. Similarly, anti-PT (APT) symmetry, characterized by $\{H,PT\}=0$, also undergoes spontaneous symmetry breaking during the APT phase transition, with EPs emerging [11, 12].
@@
Over the past decade, PT symmetry [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and APT symmetry [26, 27, 28, 29, 30, 31, 32, 33] have been extensively studied in fields like photonics, acoustics, and ultracold atoms, yielding notable experimental and theoretical progress. PT symmetry in physical systems typically requires linear gain and loss [3, 4, 26, 34]. However, implementing PT symmetry in quantum systems is challenging due to Langevin noise [35, 36, 37]. Two approaches to address this are passive schemes based on loss modulation and Hamiltonian dilation [37, 38, 39, 40, 41, 42]. Conversely, APT symmetry can be achieved without gain and loss [29, 43, 44, 45]. Theoretical research suggests that APT symmetry in a two-mode bosonic system, free from linear gain and loss, can avoid Langevin noise and be implemented in nonlinear wave-mixing processes [45]. Recent studies on quantum entanglement [46, 47] have shown unique PT and APT symmetry properties in two-qubit systems [48] and linear optical systems [49]. However, no physical systems simultaneously displaying both PT and APT symmetry have been examined.
@@
In this letter, we constructs a phase-sensitive non-Hermitian wave mixing model with inherent single-mode APT symmetry and achieves two-mode quadrature-PT symmetry by modulating the pump field phase, an aspect unexplored in prior research. Classical-nonclassical polarity (CNP) demonstrates the unique impact of APT symmetry on a system’s nonclassical properties, significantly altering single-mode squeezing and two-body Gaussian entanglement dynamics as the system traverses EP. The inseparability criterion results provide insights into the relationship between two-mode quadrature-entanglement and the pump field phase, revealing the influence of single-mode APT and two-mode quadrature-PT symmetries on entanglement. Specifically, when $\cos\phi=0$, the system exhibits both single-mode APT and two-mode quadrature-PT symmetries, with two-mode quadrature-entanglement showing sudden death in the two-mode quadrature-PT-broken region and EP, and periodic revival in the two-mode quadrature-PT-symmetric region. Conversely, when $\cos\phi=\pm 1$, the phase transition from APT-symmetric to APT-broken regions corresponds with the emergence of quantum entanglement from nonexistence, indicating a classical to quantum transition. This work further investigates the unique non-Hermitian properties of nonlinear systems, exploring the relationship between two-mode quantum entanglement and the pump field phase, elucidating the entanglement generation and evolution process in the model’s nonlinear system, offering new perspectives for quantum information processing.
@@
Figure 1: (a) A non-Hermitian system includes a four-wave mixing process and a phase-controllable pump source. Here, $\chi$ denotes the nonlinear susceptibility, $\varphi$ is the pump field phase, $\theta$ is the angle between the pump and signal (idler) fields, and $L$ is the total length of the nonlinear medium along the pump field’s optical axis. (b) APT phase transition process of the eigenvalues $\lambda_{\pm}$.
@@
Classical-Nonclassical Polarity.—To investigate the intrinsic connection between non-classical properties and the non-Hermitian nature of this phase-sensitive model, we first study Classical-Nonclassical Polarity (CNP) [50] in our model. Using the covariance matrix $\gamma_{si}$ (see supplementary materials for more detail), the single-mode and two-mode CNPs can be written as [50]
@@
$P_{1}$ indicates the single-mode squeezing property of the system, with the signal and idler fields sharing identical single-mode squeezing characteristics. $P_{2}$ represents the two-mode Gaussian entanglement. CNP greater than 0 signifies a nonclassical state, CNP of 0 marks the boundary between classical and nonclassical states, and CNP less than 0 denotes a classical state. The absolute value of the CNP measures the distance from the classical-nonclassical boundary.
In here, the nonclassical properties measured by CNP exhibit two key characteristics. First, as indicated in Eq. (5a) and (5b), the input coherent state’s intensity does not affect the nonclassical properties measured by CNP. Second, the pump field’s phase does not influence these properties, implying that single-mode squeezing and two-mode Gaussian entanglement remain unaffected by the pump field phase. The nonclassical properties mainly depend on the system parameters $\kappa$ and $\mathrm{\Delta}$, and the evolution distance $\kappa L$. Thus, discussing the CNP connection between single-mode APT symmetry and two-mode quadrature-PT symmetry here does not violate the condition $\cos\phi=0$.
@@
Fig.3 illustrates the evolution of type-I/type-II entanglement across different phases $\phi$. When $\cos\phi=0$ (e.g., $\phi=1.5\pi$ or $0.5\pi$), the system exhibits both single-mode APT symmetry and two-mode quadrature-PT symmetry, as indicated by the dashed gray line. Comparing fig.3 ($\mathrm{a}_{1}$), ($b_{1}$), and ($c_{1}$) with ($a_{2}$), ($b_{2}$), and ($c_{2}$), reveals that type-I entanglement $E_{1}$ and type-II entanglement $E_{2}$ differ only by a phase of $\pi$. Thus, the analysis focuses on type-I entanglement $E_{1}$. Fig.3($a_{1}$) demonstrates that in the single-mode APT-symmetric region ($\delta=0.8$) and when $\pi<\phi<2\pi$, the system generates entanglement, which is maintained within a narrow phase range. Outside this range, entanglement appears briefly before quickly disappearing. For $0<\phi<\pi$, no entanglement is generated. Fig.3($c_{1}$) shows that in the single-mode APT-broken region ($\delta=1.2$), system entanglement evolves oscillatory with a period $T=\kappa\pi/\lambda$, and the entanglement degree can be adjusted by varying $\phi$. Notably, Fig.3($b_{1}$) indicates that at the EP ($\delta=1$), tuning $\phi$ close to 0 maintains entanglement over a broader phase range than in the APT-symmetric region. This highlights the significant role of EP control in non-Hermitian systems and the importance of $\phi$ tuning in generating entanglement in this model.
@@
Conclusion.—In summary, unlike previous studies, this study introduces a phase-sensitive non-Hermitian four-wave mixing model, wherein the single-mode effective Hamiltonian satisfies APT symmetry. Adjustment of the pump field phase enables the two-mode effective Hamiltonian to exhibit two-mode quadrature-PT symmetry. Analysis of the system’s non-classical characteristics utilizing CNP and the inseparability criterion reveals that single-mode squeezing and two-mode Gaussian entanglement measured by CNP are phase-independent, while two-mode quadrature-entanglement measured by the inseparability criterion is phase-dependent. This distinction arises because CNP measures overall non-classicality, whereas the inseparability criterion evaluates quadrature component entanglement.
Theoretical calculations demonstrate that when $\cos\phi=0$, the system exhibits single-mode APT symmetry and two-mode quadrature-PT symmetry. Under these conditions, quantum entanglement experiences sudden death in the two-mode quadrature-PT-broken region and EP but exhibits periodic revival in the PT-symmetric region. When $\cos\phi=\pm 1$, the system transitions from APT-symmetric through EP to APT-broken regions, with quantum entanglement emerging, signifying a shift from classical to quantum behavior. This research explores the unique non-Hermitian properties of nonlinear systems, elucidates the relationship between two-mode quantum entanglement and the pump phase, and enhances understanding of entanglement generation and evolution, offering novel insights for quantum information processing.